OpenStudy (anonymous):

What is the slope of the polar curve r=cos(2 theta) at theta =0 and theta=pi/2. (This is Calculus II)

OpenStudy (amistre64):

dr/d,theta

OpenStudy (amistre64):

it may be simpler in this case to just convert it to xy coords

OpenStudy (amistre64):

parametric xy coords that is

OpenStudy (amistre64):

cos(2t) = cos^2(t) - sin^2(t) = 2cos^2(t) - 1 x = r cos(t); y=sin(t); r^2 = x^2 + y^2 those might play into a parametric solution